Project

General

Profile

2015-03-10 physics coordinator report » History » Version 15

Rogers, Chris, 10 March 2015 14:33

1 1 Rogers, Chris
h1. physics coordinator report
2
3
* Step I papers:-
4
** EMR data paper- VC talk, analysis completed and now in "write up" phase
5
** PID paper:- hung up on reconstruction issues; fixed unpacking has been run
6
   but ADC response of the KL now looks different. Discussion Monday pm
7 3 Rogers, Chris
** (Polarisation analysis - held up by dodgy EMR code)
8 1 Rogers, Chris
* Cooling demo paper:
9
** Secondary absorber did not present sufficient material to the beamline to
10
   shield the tracker, review of secondary absorber position which has caused
11
   delay to the paper.
12
** Tried for a revision this week for review, it didn't work out, hope to have a
13
   revision by next week
14
* Step IV papers:
15
** Status at http://micewww.pp.rl.ac.uk/analysis/analysis_dashboard.html
16 5 Rogers, Chris
** Trying to bring material physics people "out of the woodwork"
17 1 Rogers, Chris
* Rogers is re-establishing selection/weighting algorithms (dead since 2007)
18 6 Rogers, Chris
** aim to drive an analysis of "physics block challenge" before MPB
19
** feed into review of data rate issues, etc.
20 1 Rogers, Chris
* Field mapping
21
** Blackmore/Cobb have a good analysis of the field axis in the FC, which does not 
22
   produce the "tails" ruining the fitting, hopeful that SS will follow in time
23
   for alignment beginning of April 
24 7 Rogers, Chris
* March/April running
25
** Rogers will be physics shifter on March 21st
26
** Trying to establish physics shifters on March 28th and following
27 1 Rogers, Chris
* June running
28
** Planning note in preparation, due Friday
29
** Discuss/finalise note at a meeting Thursday 19th
30 8 Rogers, Chris
** Tracker will not be able to get plan in time for the meeting
31 10 Rogers, Chris
** Beamline (hardware) request an activation run
32 14 Rogers, Chris
** EMR request one week data and additionally some cosmics
33 8 Rogers, Chris
** C+M request one week
34 15 Rogers, Chris
** TOF request calibration (1 or 2 shifts)
35 10 Rogers, Chris
** KL request cosmic data taking
36 13 Rogers, Chris
** Ckov request 1 shift
37 10 Rogers, Chris
** Beamline (optics) not sure
38 11 Rogers, Chris
** Cooling channel first data may at some point come through, not sure